# Learning control variables and instruments for causal analysis in observational data

### v1

@@ Abstract
This study introduces a data-driven, machine learning-based method to detect suitable control variables and instruments for assessing the causal effect of a treatment on an outcome in observational data, if they exist. Our approach tests the joint existence of instruments, which are associated with the treatment but not directly with the outcome (at least conditional on observables), and suitable control variables, conditional on which the treatment is exogenous, and learns the partition of instruments and control variables from the observed data. The detection of sets of instruments and control variables relies on the condition that proper instruments are conditionally independent of the outcome given the treatment and suitable control variables. We establish the consistency of our method for detecting control variables and instruments under certain regularity conditions, investigate the finite sample performance through a simulation study, and provide an empirical application to labor market data from the Job Corps study.
@@
Addresses for correspondence: Nicolas Apfel, Murray Building (B58), Highfield Campus, University of Southampton, SO17 1TR Southampton, United Kingdom; n.apfel@soton.ac.uk. Julia Hatamyar, Research Fellow Centre for Health Economics, Heslington, York, YO10 5DD, United Kingdom; julia.hatamyar@york.ac.uk. Martin Huber, University of Fribourg, Bd. de Pérolles 90, 1700 Fribourg, Switzerland; martin.huber@unifr.ch. Jannis Kueck, Heinrich Heine University Düsseldorf, Universitätsstr. 1, 40225 Düsseldorf, Germany; kueck@dice.hhu.de.
@@ 1 Introduction
Our discussion of the test focuses on conditional mean (rather than full) independence of the instrument, which implies the identification of average treatment effects (ATE). When assuming a limited set of observed variables (relative to the sample size), we employ regression for both selecting the candidate instruments in the first step and testing the conditional mean independence of the instrument in the second step. More concisely, testing is based on the mean squared difference in outcome prediction when regressing the outcome (1) on the treatment, the control variables, and the candidate instrument and (2) on the treatment and the control variables (but not the candidate instrument). This approach is closely related to the mean squared difference test based on a quadratic score function in Huber and Kueck (2022), but applies is sequentially to all candidate instruments. We demonstrate that our method is consistent for detecting covariates and instruments under certain regularity conditions. In a simulation study, we find support for this consistency result. However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset rather with more than just several 1000 observations.
As an empirical illustration, we apply our method to data from an experimental study on Job Corps, a significant U.S. education program financed by the Department of Labor, targeting disadvantaged individuals aged 16 to 24, which has been previously analyzed in studies by Schochet et al. (2001) and Schochet et al. (2008). In this dataset, assignment to the program was randomized, making it a plausible candidate instrument for actual training participation in the first year of Job Corps, the treatment under investigation, if any direct effects of mere assignment on labor market outcomes like wage can be ruled out. Indeed, our approach selects random assignment as valid instrument. The method also suggests that training participation is exogenous conditional on a set of 28 pre-assignment covariates, with the p-value of the test statistic amounting to 76%.
@@
Our paper contributes to a growing literature on testing identifying assumptions for causal inference. For instance, de Luna and Johansson (2014) and Black et al. (2015) make use of the same conditional independence of the instrument as considered here to test the selection-on-observables assumption for the treatment when assuming a valid instrument, based on matching or regression estimators. [^1] Angrist (2004), Brinch et al. (2017), and Huber (2013) consider related tests when instrument validity is holds unconditionally without controlling for covariates, in order to test the unconditional independence of the treatment and potential outcomes. Bertanha and Imbens (2015) consider the fuzzy regression discontinuity design (RDD), where instrument validity holds at a cutoff of a running variable which discontinuously affects treatment assignment. In contrast, Huber and Kueck (2022) jointly test both instrument validity and selection-on-observables assumptions with pre-defined instruments and covariates using machine learning approaches which permit for high dimensional covariates. This is conceptually closest to our approach, but one important difference is that the method of Huber and Kueck (2022) requires specifying the sets of supposed instruments and covariates to be used for testing, whereas in this paper, these sets of covariates and instruments are learned from the data and may therefore be a priori unknown.
@@
Beyond that, we provide a new orthorgonalized quadratic score which improves on the quadratic score function used in Huber and Kueck (2022). This new score is of more general interest since it can be applied in many other contexts than ours as well. One example is the study by Parikh et al. (2024), who compare experimental and non-experimental treatment effect estimates and use an indicator for distinguishing between experimental and non-experimental evaluation designs, which has a comparable role as the instrument in our paper. Imposing the external validity of the experiment permits assessing the selection-on-observables assumption in the non-experimental design. Conversely, imposing the selection-on-observables assumption in the non-experimental design permits assessing the external validity of the experiment, which is equivalent to testing IV validity in our context. Unlike our paper, the authors do not consider testing both assumptions jointly. Also Angrist and Rokkanen (2015) use the same conditional independence of the instrument as considered here to test instrument validity within the framework of the sharp regression discontinuity design (RDD), where the selection-on-observables for the treatment holds by design, because it is a deterministic function of a cutoff in a running variable. This allows for testing whether the running variable is a valid instrument, i.e., whether it is not associated with the outcome conditional on the treatment. [^2] If the running variable is a valid instrument, causal effects can also be identified away from the cutoff, which is otherwise not feasible due to the lack of common support in the treatment across different values of the running variable. Again, our approach differs in that it tests the selection-on-observables and instrument validity assumptions jointly (rather than the latter conditional on the former).
@@ 2 Identifying assumptions and testable conditional independence
$\mathcal{D}$ and $\mathcal{Y}$ denote the support of $D$ and $Y$, respectively. The first line of Assumption 1 rules out a causal effect of outcome $Y$ on $D$, $X$, or $Z$ and of treatment $D$ on $X$ or $Z$. However, it allows for the possibility of both $X$ and $Z$ affecting $D$, $Y$, or even each other. This assumption aligns with the directed acyclic graph (DAG, see e.g. Pearl (2000)) presented in Figure 1, where causal relationships between variables are indicated by arrows: $Z$ and $X$ affect $D$, and $D$ and $X$ affect $Y$. Additionally, $X$ may influence $Z$ or vice versa, denoted by the bidirectional causal arrow. This is in line with the conventional practice of measuring covariates and instruments before treatment assignment, eliminating the potential for reverse causality between $D$ and $Y$ and the pre-treatment variables $X$ and $Z$. The DAG also features unobserved terms $U$ and $V$ that affect $Y$ and $D$, respectively, with dashed arrows denoting the unobservable nature of these effects. The second line of Assumption 1 enforces causal faithfulness, ensuring that only variables which are d-separated in the sense of Pearl (1988), i.e. not associated with each other via some causal paths (possibly conditional on other variables) are statistically independent (or conditionally independent).444 d-separation relies on blocking causal paths between variables. Formally, a path between two (sets of) variables $A$ and $B$ is blocked when conditioning on a (set of) control variable(s) $C$ if 1. the path between $A$ and $B$ is a causal chain, implying that $A\rightarrow M\rightarrow B$ or $A\leftarrow M\leftarrow B$, or a confounding association, implying that $A\leftarrow M\rightarrow B$, and variable (set) $M$ is among control variables $C$ (i.e. controlled for), 2. the path between $A$ and $B$ contains a collider, implying that $A\rightarrow S\leftarrow B$, and variable (set) $S$ or any variable (set) causally affected by $S$ is not among control variables $C$ (i.e. not controlled for). Based on this definition of blocking, the d-separation criterion states that $A$ and $B$ are d-separated when conditioning on control variable(s) $C$ if and only if $C$ blocks all paths between $D$ and $Y$. While d-separation is generally a sufficient condition for the (conditional) independence of two variables, it is a necessary condition under causal faithfulness.
@@
[…]
@@
[…]
@@
[…]
@@ 3 Testing based on double machine learning
$W=(Y,D,X,Z,\zeta)$ are random variables and $\eta=(\eta_{1},\eta_{2})$ are the so-called nuisance parameters, whose true values correspond to $\eta_{0,1}(W)=\mu(D,X,Z)$ and $\eta_{0,2}(W)=m(D,X)$. We note that the independent mean-zero random variable $\zeta$ in (3.3) is added to avoid a degenerate distribution of the estimator under $H_{0}$, a common problem in specification tests, see e.g. Hong and White (1995) and Wooldridge (1992). A disadvantage of testing based on the score function in equation (3.3) is the requirement to choose a random term $\zeta$, as the optimal selection of $\zeta$ in a given dataset is generally unknown. Further, while the estimator based on equation (3.3) is asymptotically normally distributed under the null hypothesis, as demonstrated in Huber and Kueck (2022), this is generally not the case under the alternative hypothesis. For this reason, we subsequently propose a new testing approach that is based on a refined score function that does not require user-selected random terms and entails a test statistic that is normally distributed under both the null and alternative hypotheses. [^6] We note that an alternative testing approach that could be applied in our context is suggested by Kook and Lundborg (2024)
@@
with $\theta_{0}=\mathbb{E}[\left(\mu(D,X,1)-\mu(D,X,0)\right)^{2}]+\mathbb{E}[\mu(D,X,1)-\mu(D,X,0)]$. Testing based on (3.4) corresponds to an aggregate $L_{2}$-type measure that can be used to test violations across values of $D$, $X$, and $Z$, which is common in specification tests based on nonparametric regression, see e.g. Racine (1997), Racine et al. (2006), Hong and White (1995) and Wooldridge (1992). In addition to the squared difference in conditional mean outcomes that we already encountered in the previous section, the score function notably contains a term in which the difference in conditional mean outcomes is multiplied with a difference in expressions obtained by inverse probability weighting (IPW) with the instrument propensity score. In fact, our new score above combines the orthogonalized squared score in (3.3) with the popular doubly robust score, as also considered for testing in Huber and Kueck (2022). Just as the squared difference in conditional mean outcomes, the score function $\tilde{\psi}$ is zero in expectations, $\mathbb{E}[\tilde{\psi}(W,\theta_{0},\eta_{0})]=0$, under the null hypothesis that $\theta_{0}=0$, which follows from iterated expectations, and satisfies the Neyman orthogonality property (see Appendix A.1).
@@
$+\sum_{l=1}^{L}\frac{(Y-\mu(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}-\theta.$
This score has a variance that is bounded away from zero, is zero in expectation under the null hypothesis and is Neyman-orthogonal, as formally shown in Appendix A.2. We may construct cross-fitted estimators of $\theta_{0}$ based on the score function (3.5). It is worth noting that the corresponding target parameter in (3.5) is given by
$\mathbb{E}\left[\sum_{l=1}^{L}(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))^{2}+(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right],$
@@
When implementing the testing approach, we assume an i.i.d. sample of size $n$, in which $i$ denotes the index of an observation and $W_{i}=(Y_{i},D_{i},X_{i},Z_{i})$ are the variable values of observation $i$ in that sample, with $i$in $\{1,2,..,n\}$. When testing, we apply cross-fitting as for instance discussed in Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey and Robins (2018) to avoid over-fitting due to a correlation of the estimation of the nuisance parameters and $\theta_{0}$. Therefore, we split the data into $K$ subsamples of size $N=n/K$. The cross-fitted estimator is given by
$\hat{\theta}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{N,k}[\psi_{k}(W_{i},0,\hat{\eta})],$ (3.6)
with $\hat{\eta}=(\hat{\mu},\hat{m},\hat{p}_{1},\dots,\hat{p}_{L})$. Under the regularity conditions in Assumption 6, the estimator $\hat{\theta}$ is asymptotically normal and $\sqrt{n}$-consistent, as stated in Theorem 1. The proof is provided in Appendix A.3.
@@
The following assumption needs to hold for all $n\geq 3$, $\mathbb{P}\in\mathcal{P}$ and $q>2$: (i) $\|Y\|_{\mathbb{P},q}<C$ and $\mathbb{E}[\varepsilon^{2}1(Z\in Z_{l})]>c$ (ii) Given a random subset $I$ of $[n]$ of size $N=n/K$, the nuisance parameter estimator $\hat{\eta}_{0}=\hat{\eta}_{0}((W_{i})_{i\in I^{c}})$ obeys $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2q}\leq C$, $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},4}\leq\delta_{N}$, and $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2}\leq\delta_{N}^{1/2}N^{-1/4}$ with $\mathbb{P}$-probability not less than $1-o(1)$.
@@
$\sqrt{n}\sigma^{-1}\hat{\theta}\rightsquigarrow N(\theta_{0},1),$ (3.7)

### v2

@@ Abstract
This study introduces a data-driven, machine learning-based method to detect suitable control variables and instruments for assessing the causal effect of a treatment on an outcome in observational data. Our approach tests the joint existence of instruments, which are associated with the treatment but not directly with the outcome (at least conditional on observables), and suitable control variables, conditional on which the treatment is exogenous, and learns the partition of instruments and control variables from the observed data. The detection of sets of instruments and control variables relies on the condition that proper instruments are conditionally independent of the outcome given the treatment and suitable control variables. We establish the consistency of our method for detecting control variables and instruments under certain regularity conditions, investigate the finite sample performance through a simulation study, and provide an empirical application to health data from the Oregon Health Insurance Experiment.
@@
We have benefited from comments by Niels Richard Hansen and Leonard Henckel. Addresses for correspondence: Nicolas Apfel, Department of Economics Faculty of Economics and Statistics, University of Innsbruck, Universitaetsstrasse 15, 6020 Innsbruck, nicolas.apfel@uibk.ac.at. Julia Hatamyar, Research Fellow Centre for Health Economics, Heslington, York, YO10 5DD, United Kingdom; julia.hatamyar@york.ac.uk. Martin Huber, University of Fribourg, Bd. de Pérolles 90, 1700 Fribourg, Switzerland; martin.huber@unifr.ch. Jannis Kueck, Heinrich Heine University Düsseldorf, Universitätsstr. 1, 40225 Düsseldorf, Germany; kueck@dice.hhu.de.
@@ 1 Introduction
Our discussion of the test focuses on conditional mean (rather than full) independence of the instrument, which implies the identification of average treatment effects (ATE). When assuming a limited set of observed variables (relative to the sample size), we employ regression for both selecting the candidate instruments in the first step and testing the conditional mean independence of the instrument in the second step. More concisely, testing is based on the mean squared difference in outcome prediction when regressing the outcome (1) on the treatment, the control variables, and the candidate instrument and (2) on the treatment and the control variables (but not the candidate instrument). This approach builds on the mean squared difference test based on a quadratic score function in Huber and Kueck (2022), but applies it sequentially across all candidate instruments. Unlike the original test, our procedure reverses the hypothesis structure: the null hypothesis states that identification fails, while the alternative hypothesis indicates successful identification. We demonstrate that our method is consistent for detecting covariates and instruments under certain regularity conditions. In a simulation study with ten covariates, we find support for this consistency result. However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset.
As an empirical illustration, we apply our method to health data from the Oregon Health Insurance Experiment, previously analysed by Finkelstein et al. (2012). In this experiment, low-income adults were randomly assigned the opportunity to apply for Medicaid, a public health insurance program in the United States. The random assignment provides a plausible instrumental variable for actual Medicaid enrolment - the treatment of interest - provided that assignment itself has no direct effect on health outcomes such as doctor visits. Our approach indeed selects random assignment as a valid instrument, together with several other variables, and indicates that Medicaid enrolment is exogenous, conditional on a rich set of more than 200 pre-assignment covariates capturing a wide range of socio-economic and health-related characteristics.
@@
Our paper contributes to a growing literature on testing identifying assumptions for causal inference. For instance, de Luna and Johansson (2014) and Black et al. (2015) make use of the same conditional independence of the instrument as considered here to test the selection-on-observables assumption for the treatment when assuming a valid instrument, based on matching or regression estimators. [^1] See also the related test by Chen et al. (2018), which (in contrast to other methods) requires symmetrically distributed error terms. Furthermore, Angrist (2004), Brinch et al. (2017), and Huber (2013) assume instrument validity to hold unconditionally without controlling for covariates, in order to test the unconditional independence of the treatment and potential outcomes. Bertanha and Imbens (2015) consider the fuzzy regression discontinuity design (RDD), where instrument validity holds at a cutoff of a running variable which discontinuously affects treatment assignment. In contrast, Huber and Kueck (2022) jointly test both instrument validity and selection-on-observables assumptions with pre-defined instruments and covariates using machine learning approaches which permit for high-dimensional covariates. [^2] A joint test is also proposed by Angrist et al. (2017), however, for parametric regression models with low dimensional covariates. This is conceptually closest to our approach, but one important difference is that the method of Huber and Kueck (2022) requires specifying the sets of supposed instruments and covariates to be used for testing, whereas in this paper, these sets of covariates and instruments are learned from the data and may therefore be a priori unknown.
@@
Beyond that, we provide a new orthogonalized quadratic score, which improves upon the quadratic score function used in Huber and Kueck (2022). This new score is of broader interest since it can be applied in many other contexts beyond ours. One example is the study by Parikh et al. (2024), who compare experimental and non-experimental treatment effect estimates and use an indicator to distinguish between experimental and non-experimental evaluation designs, which plays a role comparable to the instrument in our paper. Imposing the external validity of the experiment allows for assessing the selection-on-observables assumption in the non-experimental design. Conversely, imposing the selection-on-observables assumption in the non-experimental design allows for assessing the external validity of the experiment, which is equivalent to testing IV validity in our context. Unlike our paper, the authors do not consider testing both assumptions jointly. Similarly, Angrist and Rokkanen (2015) employ the same conditional independence of the instrument as considered here to test instrument validity within the framework of the sharp regression discontinuity design (RDD), where selection-on-observables for the treatment holds by design, as it is a deterministic function of a cutoff in a running variable. This allows testing whether the running variable is a valid instrument, i.e., whether it is not associated with the outcome conditional on the treatment. [^3] If the running variable is a valid instrument, causal effects can also be identified away from the cutoff, which is otherwise not feasible due to the lack of common support in the treatment across different values of the running variable.
Relatedly, one strand of the statistics literature imposes the selection-on-observables assumption in observational studies and exploits the conditional independence condition to identify subsets of covariates that are sufficient for identification, implying that the remaining covariates satisfy instrument validity; see, for instance, De Luna et al. (2011) and VanderWeele and Shpitser (2011). Considering subsets of covariate information may also allow for more efficient causal effect estimation. For example, Christgau and Hansen (2024) propose a deep learning-based estimator that learns efficient covariate representations from unstructured data and is asymptotically normal under certain conditions. Again, our approach differs from these studies in that it tests the selection-on-observables and instrument validity assumptions jointly, rather than testing the latter conditional on the former. Closer to our framework, Entner et al. (2013) do not pre-impose the selection-on-observables assumption when searching for a sufficient set of covariates to control for based on conditional independence and consider a parametric modelling approach for testing, however, without providing asymptotic guarantees. Here, we suggest a machine learning-based procedure that accommodates nonparametric models and potentially high-dimensional covariates, while having desirable asymptotic properties under certain regularity conditions.
@@ 2 Identifying assumptions and testable conditional independence
$\mathcal{D}$ and $\mathcal{Y}$ denote the support of $D$ and $Y$, respectively. The first line of Assumption 1 rules out a causal effect of outcome $Y$ on $D$, $X$, or $Z$ and of treatment $D$ on $X$ or $Z$. However, it allows for the possibility of both $X$ and $Z$ affecting $D$, $Y$, or even each other. This assumption aligns with the directed acyclic graph (DAG, see e.g. Pearl (2000)) presented in Figure 1, where causal relationships between variables are indicated by arrows: $Z$ and $X$ affect $D$, and $D$ and $X$ affect $Y$. Additionally, $X$ may influence $Z$ or vice versa, denoted by the bidirectional causal arrow. This is in line with the conventional practice of measuring covariates and instruments before treatment assignment, eliminating the potential for reverse causality between $D$ and $Y$ and the pre-treatment variables $X$ and $Z$. The DAG also features unobserved terms $U$ and $V$ that affect $Y$ and $D$, respectively, with dashed arrows denoting the unobservable nature of these effects. The second line of Assumption 1 enforces causal faithfulness, ensuring that only variables which are d-separated in the sense of Pearl (1988), i.e. not associated with each other via some causal paths (possibly conditional on other variables) are statistically independent (or conditionally independent).555d-separation relies on blocking causal paths between variables. Formally, a path between two (sets of) variables $A$ and $B$ is blocked when conditioning on a (set of) control variable(s) $C$ if 1. the path between $A$ and $B$ is a causal chain, implying that $A\rightarrow M\rightarrow B$ or $A\leftarrow M\leftarrow B$, or a confounding association, implying that $A\leftarrow M\rightarrow B$, and variable (set) $M$ is among control variables $C$ (i.e. controlled for), 2. the path between $A$ and $B$ contains a collider, implying that $A\rightarrow S\leftarrow B$, and variable (set) $S$ or any variable (set) causally affected by $S$ is not among control variables $C$ (i.e. not controlled for). Based on this definition of blocking, the d-separation criterion states that $A$ and $B$ are d-separated when conditioning on control variable(s) $C$ if and only if $C$ blocks all paths between $D$ and $Y$. While d-separation is generally a sufficient condition for the (conditional) independence of two variables, it is a necessary condition under causal faithfulness.
@@
$D\not\!\perp\!\!\!\perp Z|X,$
where $\not\!\perp\!\!\!\perp$ denotes statistical dependence. Together with Assumption 1, which rules out effects of $D$ on $Z$, Assumption 3 either implies that $Z$ causally affects $D$, which is known as first stage effect in the IV literature, or that some (unobserved) characteristics jointly affect $Z$ and $D$ given $X$. This assumption is satisfied in Figure 1, where $Z$ affects $D$.
@@
where $\perp\!\!\!\perp$} denotes statistical independence. Assumption 4 implies that conditional on covariates $X$, there exist no unobserved confounders jointly affecting outcome $Y$ and treatment $D$.
@@
Assumption 5 rules out unobserved confounders jointly affecting $Y$ and $Z$ when controlling for $X$. This is analogous to Assumption 4, but now concerning $Z$ rather than $D$. Furthermore, by assuming that the potential outcome is solely a function of $d$ (and not $z$), Assumption 5 also implies that the instrument does not have a direct impact on the outcome, except through its influence on the treatment, conditional on $X$. By this exclusion restriction, it holds that conditional on $X$, $Y(d,z)=Y(d,z^{\prime})=Y(d)$ for any instrument values $z$ and $z^{\prime}$. Otherwise, Assumption 5 would be violated, because it would follow that $Y(d)=Y(d,Z)$ and $Y(d,Z)\not\!\perp\!\!\!\perp Z|X$.
@@ 3 Testing based on double machine learning
$W=(Y,D,X,Z,\zeta)$ are random variables and $\eta=(\eta_{1},\eta_{2})$ are the so-called nuisance parameters, whose true values correspond to $\eta_{0,1}(W)=\mu(D,X,Z)$ and $\eta_{0,2}(W)=m(D,X)$. We note that the independent mean-zero random variable $\zeta$ in (3.3) is added to avoid a degenerate distribution of the estimator under $H_{0}$, a common problem in specification tests, see e.g. Hong and White (1995) and Wooldridge (1992). A disadvantage of testing based on the score function in equation (3.3) is the requirement to choose a random term $\zeta$, as the optimal selection of $\zeta$ in a given dataset is generally unknown. Further, while the estimator based on equation (3.3) is asymptotically normally distributed under the null hypothesis, as demonstrated in Huber and Kueck (2022), this is generally not the case under the alternative hypothesis. For this reason, we subsequently propose a new testing approach that is based on a refined score function that does not require user-selected random terms and entails a test statistic that is normally distributed under both the null and alternative hypotheses. [^7] We note that an alternative testing approach for conditional independence satisfying Neyman orthogonality that can be combined with machine learning and applied in our context is suggested by Lundborg et al. (2024), see also Kook and Lundborg (2024).
@@
with $\theta_{0}=\mathbb{E}[\left(\mu(D,X,1)-\mu(D,X,0)\right)^{2}]+\mathbb{E}[\mu(D,X,1)-\mu(D,X,0)]$. Testing based on (3.4) corresponds to an aggregate $L_{2}$-type measure that can be used to test violations across values of $D$, $X$, and $Z$, which is common in specification tests based on nonparametric regression, see e.g. Racine (1997), Racine et al. (2006), Hong and White (1995) and Wooldridge (1992). In addition to the squared difference in conditional mean outcomes that we already encountered, the score function notably contains a term in which the difference in conditional mean outcomes is multiplied with a difference in expressions obtained by inverse probability weighting (IPW) with the instrument propensity score. In fact, our new score above combines the orthogonalized squared score in (3.3) with the popular doubly robust score, as also considered for testing in Huber and Kueck (2022). Just as the squared difference in conditional mean outcomes, the score function $\tilde{\psi}$ is zero in expectations, $\mathbb{E}[\tilde{\psi}(W,\theta_{0},\eta_{0})]=0$, under the null hypothesis that $\theta_{0}=0$, which follows from iterated expectations, and satisfies the Neyman orthogonality property (see Appendix A.1).
@@
$+\sum_{l=1}^{L}\left(\frac{(Y-\mu(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right).$
This score has a variance that is bounded away from zero, is zero in expectation under the null hypothesis, $\theta_{0}=0$, and is Neyman-orthogonal, as formally shown in Appendix A.2. We may construct cross-fitted estimators of $\theta_{0}$ based on the score function (3.5). It is worth noting that the corresponding target parameter in (3.5) is given by
$\mathbb{E}\left[\sum_{l=1}^{L}[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))^{2}+(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))]\right],$
@@
When implementing the testing approach, we assume an i.i.d. sample of size $n$, in which $i$ denotes the index of an observation and $W_{i}=(Y_{i},D_{i},X_{i},Z_{i})$ are the variable values of observation $i$ in that sample, with $i$ $\in$ $\{1,2,..,n\}$. When testing, we apply cross-fitting as for instance discussed in Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey and Robins (2018) to avoid over-fitting due to a correlation of the estimation of the nuisance parameters and $\theta_{0}$. Therefore, we split the data into $K$ subsamples of size $N=n/K$. The cross-fitted estimator is given by
[…]
with $\hat{\eta}=(\hat{\mu},\hat{p}_{1},\dots,\hat{p}_{L})$. Under the regularity conditions in Assumption 6, the estimator $\hat{\theta}$ is asymptotically normal and $\sqrt{n}$-consistent, as stated in Theorem 1. The proof is provided in Appendix A.3.
@@
Define $U=Y-\mu(D,X,Z)$. The following assumption needs to hold for all $n\geq 3$, $\mathbb{P}\in\mathcal{P}$ and $q>2$: (i) $\|Y\|_{\mathbb{P},q}<C$ and $\mathbb{E}[U^{2}1(Z\in Z_{l})]>c$ (ii) Given a random subset $I$ of $[n]$ of size $N=n/K$, the nuisance parameter estimator $\hat{\eta}_{0}=\hat{\eta}_{0}((W_{i})_{i\in I^{c}})$ obeys $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2q}\leq C$, $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},4}\leq\delta_{N}$, and $\|\hat{\eta}-\eta_{0}\|_{\mathbb{P},2}\leq\delta_{N}^{1/2}N^{-1/4}$ with $\mathbb{P}$-probability not less than 1o(1)1-o(1).
@@
$\sqrt{n}\sigma^{-1}\hat{\theta}\overset{d}{\rightarrow}N(\theta_{0},1),$ (3.7)